3. Stocks that pay a high dividend are called ____________ stocks.

  1. Growth
  2. Large Cap
  3. Income
  4. Speculative

4. The stock price will fall by the amount of the dividend on the _____________.

  1. Declaration Date
  2. Ex Dividend Date
  3. Payment Date
  4. Record Date

5. What is the value of a right if the stock (including the right) sells for $60 and it takes four rights to buy stock for $40?

a) $4                         b) $4.5                    c) $5                         d) $6.67
Answer #1

3)stocks that tends to pay high dividend is called "Income stock " since they are providing income by way high dividend.

4)correct option is "b"- Ex Dividend Date

As soon as dividend is paid ,price of a stock goes ex dividend (that is price of stock falls by amount of dividend)

5)correct option is "c"-$ 5

Benefit due to right = 60-40 = $ 20

Price per right = 20/4 =$ 5 per right